OpenStudy (anonymous):

A little help pwease :o ? A group of 45 people attended a ball game. There were twice as many children as adults in the group. Set up a system of equations that represents the numbers of adults and children who attended the game and solve the system to find the number of children who were in the group. O.o

OpenStudy (anonymous):

I know the equation is a+c=45 btw haha.

OpenStudy (anonymous):

Let # of Children=C, # of Adults=A So since total # of people=45, hence C+A=45 Also # of children= twice the # of adults Hence C=2A System of Equations i) C+A=45 ii) C=2A Now substitute ii) in i) to get the value of A and from that get the value of C

OpenStudy (anonymous):

oh alrighty haha

OpenStudy (anonymous):

5. A group of 45 people attended a ball game. There were twice as many children as adults in the group. Set up a system of equations that represents the numbers of adults and children who attended the game and solve the system to find the number of children who were in the group. (1 point) (1 pt) 15 adults, 30 children (0 pts) 30 adults; 22 children (0 pts) 22 adults; 30 children (0 pts) 30 adults, 15 children 1 /1 point